### v3

# EntroLLM: Entropy Encoded Weight Compression for Efficient Large Language Model Inference on Edge Devices

Abstract
Large Language Models (LLMs) demonstrate exceptional performance across various tasks, but their large storage and computational requirements constrain their deployment on edge devices. To address this, we propose EntroLLM, a novel compression framework that integrates mixed quantization with entropy coding to reduce storage overhead while maintaining model accuracy. Our method applies a layer-wise mixed quantization scheme – choosing between symmetric and asymmetric quantization based on individual layer weight distributions – to optimize compressibility. We then employ Huffman encoding for lossless compression of the quantized weights, significantly reducing memory bandwidth requirements. Furthermore, we introduce parallel Huffman decoding, which enables efficient retrieval of encoded weights during inference, ensuring minimal latency impact. Our experiments on edge-compatible LLMs, including smolLM-1.7B-Instruct, phi3-mini-4k-Instruct, and mistral-7B-Instruct, demonstrate that EntroLLM achieves up to $30\%$ storage reduction compared to uint8 models and up to $65\%$ storage reduction compared to uint4 models, while preserving perplexity and accuracy, on language benchmark tasks. We further show that our method enables $31.9\%$$146.6\%$ faster inference throughput on memory-bandwidth-limited edge devices, such as NVIDIA Jetson P3450, by reducing the required data movement. The proposed approach requires no additional re-training and is fully compatible with existing post-training quantization methods, making it a practical solution for edge LLMs.
Index Terms:
Large Language Models (LLMs), quantization, entropy coding, Huffman coding, parallel decoding
I Introduction
Large language models (LLMs) have demonstrated remarkable performance in various domains [1, 2], but their substantial size poses challenges for deployment, especially on resource-constrained edge devices [3]. For example, even a smaller LLM such as mistral-7B-Instruct [4] requires over $14$ GB of memory with weights encoded in $16$-bit floats (fp16), exceeding the capacity of most edge devices. As a direct result of being massively parameter-heavy, the primary bottleneck in LLM inference, particularly for generative tasks, is memory bandwidth rather than computation [5]. This means that the speed of loading and storing parameters becomes the most critical constraint [6].
Figure 1: Overall Schematic that demonstrates our edge-device inference scheme. A floating point model is first quantized using a mixed quantization scheme to maximize storage compressibility. The model after quantization (blue) is the model that we will finally use during inference. But when storing this model on edge devices (green), we compress the integer weights using entropy coding. To ensure fast inference, we come up with a parallel decoding scheme to revert to integer weights from variable length entropy-coded weights. Section III has a detailed description of the various components described here.
Figure 2: A visual explanation of the different uniform quantization grids [7] for a bit-width of $8$. $s$ is the scaling factor, $z$ the zero-point. The floating-point grid is in black, and the integer quantized grid is in blue. In our work, we use either a symmetric unsigned or an asymmetric quantization scheme on each layer based on the individual layer’s weight distribution.
Quantization, a technique that encodes the model parameters in lower precision formats, has effectively reduced the size of LLMs [8]. However, further compression without compromising performance remains challenging, especially for smaller LLMs ($\leq$ 10B parameters) designed for edge applications. To address this, we focus on reducing the storage overhead for deploying these small LLMs on low-power edge devices. We introduce an entropy-based encoding scheme to further compress these models without loss of information. Specifically, we leverage variable-length bit widths to encode the quantized weights, using techniques rooted in entropy coding to achieve lossless compression. During inference, the original weights are recovered with precision without additional storage overhead, ensuring that performance degradation is minimized to minor quantization effects. This approach, illustrated in Figure 1, employs Huffman coding due to its alignment with the Shannon entropy bounds, offering optimal compression for the quantized model parameters while maintaining the computational efficiency required for edge inference. We also use mixed asymmetric and symmetric unsigned quantization to induce better weight compression and implement a partitioning scheme for model weights so that parallel decoding is made possible during inference.
Our main contributions to this work are as follows –
  1. 1.
    Mixed Quantization Scheme: We introduce a layer-specific quantization strategy that balances storage efficiency with model performance, ensuring optimal weight distribution across layers.
  2. 2.
    Huffman Weight Encoding: We leverage entropy-based Huffman coding to further compress quantized weights, reducing storage overhead while preserving inference accuracy.
  3. 3.
    Parallel Decoding for Efficient Inference: To mitigate latency issues caused by sequential entropy decoding, we implement a parallelized decoding strategy that enables real-time processing on edge hardware.
  4. 4.
    Comprehensive Evaluation: We evaluated our approach on multiple LLMs (smolLM-1.7B-Instruct [9], phi3-mini-4k-Instruct [10], and mistral-7B-Instruct [4]) deployed on devices with limited resources, demonstrating significant storage savings and reduced inference latency.
II Related Works
Before the era of LLMs, various compression techniques were developed to enhance the efficiency of deep learning models. These techniques include pruning [11], quantization [12], knowledge distillation [13], and alternative numerical representations such as logarithmic numbers [14] and posit [15]. Although effective, most of these methods rely on full or partial retraining of the models to achieve optimal results. However, for LLMs, retraining is highly impractical, as their memory requirements during training exceed the capacity of most consumer and commercial GPUs available today. To bypass this challenge, recent research has proposed compression methods for LLMs that require minimal to no retraining.
II-A Pruning
Sun et al. [16] proposed an approach that prunes weights with the smallest magnitudes multiplied by the corresponding input activations per output. LLM-pruner [17] adopted structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the major functionality of the LLM.
II-B Quantization
Numerous weight-only post-training quantization methods have been introduced, such as GPTQ [18], AWQ [19], and SpQR [20], to enhance the memory bandwidth of LLMs. To minimize accuracy loss, these techniques typically leverage a small calibration dataset to fine-tune the quantized weights. The primary goal of these approaches is to maintain the layer-wise output activations consistent with their full-precision counterparts, allowing weight precision to be reduced to $4$ bits with minimal performance impact. Although these works focus on uniform quantization and are commonly used, others focus on non-uniform quantization based on sensitivity [21, 5]. These works focus on mitigating the effects of parsimony on more sensitive weight values, allowing more aggressive quantization beyond $8$ bits without compromising performance. Recent studies have also explored ultra-low-precision models [22, 23], such as ternary LLMs [24], which address outliers in both activations and weights. However, the applicability of these advanced quantization techniques to smaller LLMs designed for edge devices remains largely unexplored.
II-C Codebook-based Entropy Coding
Codebook-based entropy coding [25] compresses model parameters by mapping frequent symbols to shorter codes using predefined dictionaries. Although this method is effective in reducing encoding length and offers sub-1-bit compression, it is primarily suited for mixture-of-experts models and lacks a practical solution for resource-constrained edge devices. Moreover, codebook-based compression is not Shannon-rate optimal, which limits its efficiency. In contrast, our approach employs Huffman encoding, which achieves optimal compression rates, and we further enhance inference efficiency by implementing parallel Huffman decoding to meet the stringent performance demands of edge hardware inference.
III Proposed Method
Entropy coding aims to represent the data in the most compact way possible by leveraging the frequency of symbols in the source material. This approach is grounded in Shannon’s entropy, which measures the average amount of information produced by a stochastic data source. Huffman encoding [26], developed by David A. Huffman in 1952, optimally achieves this by assigning shorter codes to more frequent symbols and longer codes to less frequent ones, ensuring that the total size of the encoded data is minimized. In our approach, Huffman encoding is our main workhorse for achieving beyond-quantization on-device compression.
III-A Mixed Quantization Scheme
In this work, we aim to compress the weights further beyond post-training quantization. The weights of a neural network are normally distributed [27] after training, and after quantization, their distribution should remain unchanged. To increase compressibility from entropy coding, we need to ensure that the quantized weights are distributed in a way that minimizes entropy. To achieve this, we employ a mixed quantization scheme that quantizes the parameters of each layer with symmetric unsigned quantization (equation 1) or asymmetric quantization (equation 2), depending on the weight distribution of that layer. The different quantization schemes and how they map from floating to integer grids are explained in figure 2.
$\mathtt{W_{int}}$ $\leftarrow\left\lfloor\frac{\mathtt{W_{fp}}}{s}\right\rceil$ (1)
$\mathtt{W_{int}}$ $\leftarrow\left\lfloor\frac{(\mathtt{W_{fp}}-z)}{s}\right\rceil$ (2)
Figure 3: In the serial decoding setting, parallelization is not possible as the variable code lengths make it difficult for us to predict the start token of a symbol in our remaining encoded parameters. By keeping the parameter space’s original weight tensor packing structure intact, we can parallelize decoding as now we can assign different encoded tensors (chunks), to individual threads assigned to cores. Modern LLMs have hundreds to thousands of such weight tensors, and hence, coarse-grained parallelism is achievable. In the figure <p> represents current bits being decoded, while <unk> represents unknown encoded parameter bits.
Algorithm 1 Entropy-based LLM Compression and Inference
1:  Input: Pre-trained LLM with floating-point weights $\mathtt{W_{fp}}$
2:  Output: Compressed model for efficient edge deployment
3:  procedure Cloud Processing $(\mathcal{C})$
4:      for each layer $\mathtt{k}$ in the model $\triangleright$ Quantization
5:          if $\mathtt{W_{fp}^{k}|_{max}\times W_{fp}^{k}|_{min}\geq 0}$ then
6:              $\mathtt{W_{int}^{k}}\leftarrow\left\lfloor\frac{\mathtt{W_{fp}^{k}}}{s}\right\rceil$
7:          else
8:              $\mathtt{W_{int}^{k}}\leftarrow\left\lfloor\frac{(\mathtt{W_{fp}^{k}}-z)}{s}\right\rceil$    $\triangleright$ $z$ is zero-point, $s$ is scaling factor
9:          end if
10:      end for
11:      $F\leftarrow\mathcal{F}\{\mathtt{W_{int}^{0}},\mathtt{W_{int}^{1}},\dots,\mathtt{W_{int}^{k}},\dots\}$$\triangleright$ Compute    frequency of each quantized weight value across model
12:      $H,P\leftarrow\mathcal{H}\{F\}$$\triangleright$ Construct Huffman tree to generate    symbol lookup table and probability statistics
13:      for each layer $\mathtt{k}$ in the model do
14:          $\mathtt{W_{c}^{k}}\leftarrow H\{\mathtt{W_{int}^{k}}\}$
15:      end for
16:      Store model metadata: $H,P,\{\mathtt{W_{c}}\}^{\mathtt{k}}$     $\triangleright$ Preserving    the weight tensor packing structure $\{\mathtt{W_{c}}\}^{\mathtt{k}}$    $\triangleright$ Enables parallel decoding during inference
17:  end procedure
18:  procedure Edge Device Operations $(\mathcal{D})$
19:      Load: $H,P,\{\mathtt{W_{c}}\}^{\mathtt{k}}$
20:      Initialize: parallel decoding with $T$ threads
21:      for each thread $t\in T$ in parallel do
22:      Assign: $\{\mathtt{W_{c}}\}^{\mathtt{m}}$ to thread $t\in T$ to balance workload
23:          for each assigned layer $\mathtt{m}$ do
24:              $\mathtt{W_{int}^{m}}\leftarrow H^{-1}\{\mathtt{W_{c}^{m}}\}$
25:              Perform layer computation using $\mathtt{W_{int}^{m}}$
26:          end for
27:      end for
28:  end procedure
This dual quantization scheme ensures that each layer retains its inherent distribution, enabling the network to operate within a consistent and normalized range across layers. By carefully choosing the quantization method based on the sign characteristics of each layer, we normalize the parameter distributions, resulting in a system in which all layers, regardless of their original distribution, contribute to a final, highly Gaussian distribution for all parameters in the model. This convergence of distributions is the key to our subsequent Huffman encoding step. Using the low entropy produced by combining these balanced weight distributions, we generate an efficient Huffman tree, allowing us to encode the model weights with minimal redundancy while maintaining precision during inference. This synergy between the quantization strategy and entropy coding not only compresses the model effectively but also ensures that the performance remains unaffected by the transformation. Lines 4 to 10 in algorithm 1 describe the quantization scheme used.
(a) smolLM-1.7B-Instruct
(b) phi3-mini-4k-Instruct
(c) mistral-7B-Instruct
(d) smolLM-1.7B-Instruct
(e) phi3-mini-4k-Instruct
(f) mistral-7B-Instruct
Figure 4: Top: Parameter distribution for 8-bit quantized models. Bottom: Parameter distribution for 4-bit quantized models. Model size increases from left to right.
Property Models
smolLM-Instruct phi3-mini-4k-Instruct mistral-Instruct
Parameter Count $1.7$ Billion $3.8$ Billion $7.0$ Billion
Weight Encoding fp16 uint8 uint4 fp16 uint8 uint4 fp16 uint8 uint4
Effective Bits $16$ $\mathbf{5.92}$ $\mathbf{1.57}$ $16$ $\mathbf{5.58}$ $\mathbf{1.39}$ $16$ $\mathbf{5.84}$ $\mathbf{1.62}$
Wikitext2 (ppl.) $\downarrow$ $\mathbf{23.81}$ $23.93$ $24.14$ $\mathbf{9.03}$ $9.44$ $10.10$ $\mathbf{8.17}$ $8.24$ $8.29$
HellaSwag (acc.) $\uparrow$ 25.85% $25.55\%$ $25.30\%$ 82.2% $82.10\%$ $81.01\%$ 58.37% $58.33\%$ 58.21%
GSM8k CoT (acc.) $\uparrow$ - - - 77.37% $72.84\%$ $70.58\%$ 52.2% $48.62\%$ $45.36\%$
TABLE I: Benchmarks: Perplexity & Accuracy benchmarks for smolLM-1.7B-Instruct, phi3-mini-4k-Instruct and mistral-7B-Instruct on various language tasks
III-B Huffman weight encoding
Huffman encoding – which originates from entropy coding – is a popular data compression technique. The technique is based on the construction of a binary tree, where each symbol is assigned a score corresponding to its frequency of occurrence. This process involves recursively combining pairs of symbols into a tree structure. More frequent symbols are positioned closer to the root of the tree, which results in shorter paths and, consequently, shorter binary codes for these symbols. The length of each symbol’s binary code is proportional to its distance from the root, thereby minimizing the total length of the encoded data. This ensures that no other prefix code can offer a more efficient representation of the data, making the Huffman encoding optimal for lossless compression.
In this work, we employ Huffman encoding to compress the weights of quantized LLMs, enabling their deployment on edge devices. By applying Huffman encoding to the quantized weights of an LLM, particularly those with skewed value distributions, we achieve significant compression without sacrificing accuracy. As discussed in Section II-B, a post-training quantized LLM can match the performance of a full-precision LLM for most language generation benchmarks. Huffman coding converts the model’s weight parameters into variable-length codes, assigning shorter codes to frequent weight values, which reduces both storage and bandwidth costs. Lines 11 to 16 in algorithm 1 describe the Huffman encoding of weights post-quantization and how the encoded weights are stored to ensure on-the-fly parallel decoding during inference.
III-C Parameter space segmentation for parallel Huffman decoding
To achieve efficient real-time inference with Huffman-encoded model weights, we introduce a method for parallelizing the decoding process. Huffman decoding is inherently sequential, as it involves interpreting variable-length encoded symbols one at a time. This sequential nature increases the latency of the LLMs when deployed on edge devices, limiting their potential for real-time execution. To overcome this limitation, we propose a segmentation strategy that preserves the original structure of the encoded weights’ parameter space during Huffman encoding.
The encoded parameter space is divided into discrete segments, maintaining the inherent organization of the weight tensors. This preservation of the structure allows for each segment to be decoded independently. Each of these segments can contain an average number of symbols and can be decoded in parallel, allowing concurrent decoding across multiple threads. By preserving the weight-tensor packing structure, we ensure that the beginning and end of each encoded segment are known in advance. This enables independent decoding without the need for synchronization between threads. By doing this, we can distribute the segments across multiple threads, allowing them to be decoded in parallel. The design ensures that each thread operates on its portion of the encoded data without having to wait for others to complete, thus bypassing the serial bottleneck that typically hinders Huffman decoding.
An additional challenge arises from the fact that individual segments can exhibit skewed distributions of Huffman-encoded symbols, leading to potential imbalances in decoding times. To address this, we employ a shuffling mechanism in which multiple segments are assigned to each thread. This ensures that, on average, each thread processes a balanced workload, as the mixture of segments helps to even out any variance in decoding complexity caused by the skewed distributions of longer or shorter symbols. This approach not only accelerates the overall decoding process but also guarantees that the system is well-suited for real-time inference scenarios, where maintaining low latency is paramount. By parallelizing the decoding of Huffman-encoded weights while preserving the original parameter space structure, we achieve both efficient model restoration and the rapid inference required in time-critical applications. The schematic for parallel decoding is represented diagrammatically in figure 3. The algorithm 1 demonstrates the proposed methodology end-to-end as described in sections III-A, III-B, and III-C.
IV Experiments
We evaluated our proposed compression scheme on three edge-based LLMs: smolLM-1.7B-Instruct (1.7 billion parameters) [9], phi3-mini-4k-Instruct (3.8 billion parameters) [10] and mistral-7B-Instruct (7 billion parameters) [4]. Table I shows their baseline fp16 sizes and subsequent sizes after quantization and Huffman encoding. Figure 4 visually confirms the Gaussian distribution of the quantized weights.
IV-A Storage Reduction
Task Encoding Latency w/o Huffman Latency w/ Huffman
(sec) (sec)
pre-fill uint8 $27.10$ $\mathbf{23.17}$
token generation $0.083$ $\mathbf{0.063}$
parallel decoding - $6.66$
first token latency $\mathbf{27.18}$ $29.89$
pre-fill uint4 $9.69$ $\mathbf{8.34}$
token generation $0.062$ $\mathbf{0.025}$
parallel decoding - $1.66$
first token latency $\mathbf{9.75}$ $10.03$
TABLE II: Latency breakdown for the phi3-mini-4k model on an NVIDIA Jetson P3450 across different quantization formats (uint8 and uint4) with and without Huffman compression.
Storage reduction is achieved through a two-step process. First, the models undergo quantization. For example, the phi3-mini-4k-Instruct model’s size is reduced from $7.2$ GB to $3.6$ GB and $1.8$ GB when quantized from fp16 to uint8 and uint4, respectively. This initial step significantly reduces the memory footprint of the models.
Following quantization, Huffman compression is applied. This further reduces the effective bit lengths from $8$ and $4$ bits (to $5.58$ and $1.39$ bits, respectively, in the case of phi3-mini-4k-Instruct). Table I provides additional details for the bit widths for the other two models, smolLM-1.7B-Instruct and mistral-7B-Instruct.
The transition from $8$-bit to $4$-bit quantization increases the skewness of the distribution. This shift happens because reducing bit-width compresses the number of discrete values available to represent parameters. In the $8$ bit quantization, the $256$ symbols allow for a more granular weight representation, while the $4$-bit quantization reduces this to only $16$ symbols. This reduction causes a “bucketing” effect, where multiple nearby values are mapped to the same symbol, reducing the randomness or entropy of the distribution. The central buckets, which contain the most frequent values, dominate the quantized distribution, especially in the 4-bit quantized model.
IV-B Performance
We tested our compressed models on a range of tasks to evaluate their performance under low-precision quantization. Specifically, we considered text generation on the Wikitext2 benchmark, common sense reasoning on the 5-shot HellaSwag benchmark, and arithmetic reasoning on the 8-shot GSM8k CoT benchmark. Table I summarizes the results for three model families—smolLM-1.7B-Instruct, phi3-mini-4k-Instruct mistral-7B-Instruct – evaluated across fp16, uint8, and uint4 precision levels.
Across all models and benchmarks, the uint8 and uint4 variants exhibit only a modest performance degradation compared to the fp16 baseline. For example, on Wikitext2, perplexity remains largely stable – phi3-mini-4k-Instruct shifts from $9.03$ to $10.10$, and mistral-7B-Instruct degrades slightly from $8.17$ to $8.29$, despite operating at just $1.39$ and $1.62$ effective bits, respectively, in the uint4 format. These results highlight the resilience of the language modeling performance of our LLMs to aggressive quantization.
For HellaSwag, accuracy remains largely consistent across bit widths, especially for larger models. The model phi3-mini-4k-Instruct, for example, drops only $1.2$ percentage points between fp16 and uint4, maintaining over $81\%$ accuracy. The smaller smolLM-1.7B-Instruct shows a slightly larger relative drop, suggesting that model size plays a role in quantization robustness.
For GSM8k CoT, which is more sensitive to precision due to its dependence on intermediate reasoning steps, we observe a larger performance gap. The phi3-mini-4k-Instruct model drops from $77.37\%$ to $70.58\%$ in accuracy when moving from fp16 to uint4, while mistral-Instruct decreases from 52.2% to 45.36%. However, given the compression achieved, this drop may still be acceptable for many deployment scenarios.
IV-C Hardware Efficiency & Parallel Decoding Performance
Table II presents latency measurements for various stages of inference using the phi3-mini-4k-Instruct model, which contains approximately $3.8$ billion parameters. The evaluation was performed on an NVIDIA Jetson P3450, where we parallelized the decoding operation across four CPU threads.
NVIDIA Jetson P3450 features a quad-core ARM Cortex-A57 CPU running at $1.43$ GHz, $4$ GB LPDDR4 memory with $25.6$ GB/s bandwidth, $32$ KB L1 data cache per core, $48$ KB L1 instruction cache per core, $2$ MB shared L2 cache, and a $128$-core Maxwell GPU. The Cortex-A57 cores support NEON SIMD instructions [28], enabling efficient bit-level parallelism via packing in C++ and significantly accelerating low-precision Huffman decoding operations like those used in our uint4 and uint8 models.
Our results show that parallel decoding is highly efficient, especially under aggressive quantization. For the uint4 model, decoding takes only $1.66$ seconds, a small number compared to the $9.69$ seconds required for pre-fill and $9.75$ seconds required to generate the first output token. Even with uint8, parallel Huffman decoding is completed in $6.66$ seconds, which remains a small fraction of the overall inference time. Crucially, Huffman decoding is performed only once per sequence. As a result, its cost is amortized over the full output length, making the overhead negligible in the overall inference pipeline. This is especially significant in benchmarks like Wikitext2, HellaSwag, and GSM8k, where each input may result in dozens or hundreds of tokens. The ability to decode all quantized weights efficiently upfront enables fast, low-power inference at the edge, showing that our compressed models are not only memory and bandwidth-efficient but also latency-friendly for real-world deployments.
IV-D Latency Reduction from Huffman Coding
Table II demonstrates the performance benefits of Huffman coding on the phi3-mini-4k model running on an NVIDIA Jetson P3450. The data reveal significant throughput and latency improvements when applying Huffman coding on quantized weights of $8$-bit and $4$-bit. For the $8$-bit (uint8) format, Huffman coding reduces the effective bit width to $5.58$ bits, yielding a $14.5\%$ reduction in pre-fill time (from $27.10$ to $23.17$ seconds) and a $31.9\%$ decrease in token generation latency (from $0.083$ sec to $0.063$ sec). Even more impressive, for $4$-bit (uint4) weights, Huffman compression achieves an effective bit-width of just $1.39$ bits, resulting in a $13.3\%$ pre-fill time reduction and a remarkable $146.6\%$ improvement in token generation latency (from $0.062$ sec to $0.025$ sec). These results demonstrate that reducing the memory bandwidth requirements through Huffman coding has a dramatic impact on inference performance, particularly for token generation. Performance gains from Huffman coding are particularly pronounced in the LLM decoding phase, which is primarily memory-bandwidth limited [29, 30]. With reduced bit-width, less data needs to be transferred from memory, directly improving the latency to generate an output token. For example, when $8$-bit weights are compressed to $5.58$ bits, theoretically $30\%$ fewer data needs to be moved, approaching a 1.43×1.43$\times$1.43 × potential speedup. Our measured 1.32×1.32$\times$1.32 × improvement $\left(\frac{15.93}{12.08}=1.32\times\right)$ closely approaches this theoretical limit, with the difference attributable to bit-packing overheads. The improvements in pre-fill latency are more modest because this phase is compute-dominated rather than memory-bandwidth limited [31]. Although our approach reduces memory access with lower bit-width, it does not reduce the computational complexity since Huffman encoded weights are decoded back to their original precision ($4$ or $8$ bits) before computation. Therefore, the primary benefit during pre-fill comes from reduced data transfer from on-chip memory and IO, not from the compute operations themselves. As discussed in Section IV-C, Huffman decoding is performed once per sequence and introduces negligible overhead, especially when amortized across long outputs common in LLM workloads.
The implementation challenges of non-standard bit formats are effectively managed through optimized CUDA kernels that efficiently pack and unpack these fractional bit-width values [32]. While current Ampere GPU Tensor Cores natively support int8 and int4 operations but not arbitrary bit-widths as obtained during Huffman coding, our approach minimizes overhead by keeping the computational precision unchanged and focusing on memory transfer optimization.
V Conclusions & Discussions
In this work, we have developed a highly efficient method for compressing and deploying large language models (LLMs) on resource-constrained edge devices. Our approach combines low-bit quantization, entropy coding via Huffman compression, and a parallelized decoding strategy to achieve significant storage and bandwidth reductions – without compromising model accuracy or inference throughput. Specifically, our mixed symmetric and asymmetric quantization strategy effectively preserves the natural weight distribution across layers, optimizing the compressibility of the quantized tensors. This enables Huffman coding to reduce the average bit-width to as low as $1.39$ bits for $4$-bit weights, yielding dramatic improvements in memory efficiency. To address the inherent latency challenges of entropy decoding, we introduce a lightweight address space segmentation scheme that enables parallel Huffman decoding, ensuring that the decompression overhead remains negligible and scalable even on devices with limited compute, such as the NVIDIA Jetson P3450. Experimental results validate the efficiency of our method in the real world: we observe substantial improvements in token generation latency, especially in memory-bound inference phases such as LLM decoding. In general, our pipeline facilitates real-time, low-power inference for quantized LLMs on embedded platforms, making it suitable for on-device natural language processing in applications such as robotics, Internet-of-things, autonomous vehicles, and edge AI assistants. Our future work includes exploring adaptive entropy coding and hardware-aware optimizations to further reduce overhead and improve deployment efficiency.